OpenStudy (asdasdasd):

Please Help Me!!! Will give medals! How can polynomial functions be written when given the zero? I learn better with examples..

OpenStudy (asdasdasd):

@campbell_st please help

OpenStudy (campbell_st):

well if you know the zeros then if x = b and x = c are zeros then the factors are x - b and x - c so the polynomial can be written as P(x) =a(x - b)(x -c) the value of a is just a multiplier... e.g if x = 2 and x = -4 are zeros find the polynomial and the point (1, -10) is on the curve then P(x) = a(x-2)(x + 4) \[P(x) = a(x^2 +2x -8)\] so substitute x = 1 and P(x) = -10 to find the value of a then a = 2 so the polynomial is \[P(x) = 2(x^2 +2x - 8)\] hope it helps

OpenStudy (asdasdasd):

thanks it did.